### v2

# Word and Phrase Features in Graph Convolutional Network for Automatic Question Classification

Abstract
Effective question classification is crucial for AI-driven educational tools, enabling adaptive learning systems to categorize questions by skill area, difficulty level, and competence. This classification not only supports educational diagnostics and analytics but also enhances complex tasks like information retrieval and question answering by associating questions with relevant categories. Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in natural language, leading to suboptimal performance. To address this, we propose a novel approach leveraging graph convolutional networks, named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions. By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows the model to learn from the interconnected nature of language more effectively. Additionally, we explore the incorporation of phrase-based features to enhance classification performance on question datasets of various domains and characteristics. Our findings demonstrate that the proposed model, augmented with these features, offer a promising solution for more robust and context-aware question classification, bridging the gap between graph neural network research and practical educational applications of AI.
Index Terms:
natural language processing, graph convolutional network, question classification
11footnotetext: Equal contribution
I Introduction
Question classification is vital in providing AI-driven education and training, assisting adaptive learning systems by categorizing questions into skill area or level of difficulty and competence, and providing educational diagnosis and analytics. Accurate automatic question classification also provides headway to more complex tasks such as information retrieval and question answering, in which the association of the given question with a certain category allows for retrieval of relevant contextual knowledge to formulate compatible answers. Traditional methods, relying on word embeddings and conventional classifiers, often struggle with capturing the nuanced relationships between words and phrases, leading to suboptimal classification performance. The complexity of natural language, especially in the context of educational and domain-specific questions, requires more sophisticated approaches that can understand and leverage the inherent structure of language.
Aside from text embeddings, graphs have recently been highlighted as a way to represent unstructured text data. Questions can be naturally represented as graphs, where nodes correspond to words or phrases, and edges capture the relationships between them, such as syntactic dependencies, semantic similarities, or proximity measures. This graph-based representation aligns perfectly with the strengths of graph convolutional network (GCN), which excels at learning from structured data. GCNs are designed to operate on graph-structured data, making them ideal for processing and understanding the interconnected nature of words and phrases within questions. This capability enables GCNs to capture the underlying structure and dependencies in a question, leading to more accurate and context-aware classification decisions. GCNs offer a promising solution by modeling the relationships between words, phrases, and their syntactic and semantic connections as graphs. This approach allows for a more nuanced and context-aware classification, bridging the gap between basic AI research and practical applications in education and beyond.
Question classification poses a unique problem even among the different types of text classification, as the amount of information available in a question tends to be limited compared to document-level text counterparts. While neither tackling the task of text classification with GCNs, nor automated question classification via artificial intelligence is new, our investigation aims to determine the feasibility of adapting GCNs for question classification. Up till our work, GCNs have not been thoroughly explored in the context of question classification, with the exception of [1] and [2] which are reviewed in Section II-A. In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings.
II Related Work
II-A Automatic Question Classification
One of the earliest works on Question Classification was in [3] which made use of support vector machines (SVMs) and explored a variety of features like lexical, synctactic, and semantic features. Another reference work by [4] also discusses usage of head words and hypernyms as features for SVM and maximum entropy (ME) algorithms. They demonstrated the effectiveness of statistical learning methods in classifying questions, while rule-based classification [5] via string matching in questions have also been explored in earlier works.
Following these works, a review of automatic question classifiers from 2012 to 2017 [6] revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods. The approaches include concatenating pre-trained word embeddings to form a question embedding [7], incorporating subword information in character-based convolutional neural network (CNN) [8], and combining CNN and long short-term memory (LSTM) in a hybrid model [9]. These neural networks, effectively balancing feature extraction and sequence modeling, were shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
While there are other works which focus on question classification using neural networks, we review the two works up to date that specifically makes use of text graphs and highlight the improvements we propose with PQ-GCN. The first work [1] incorporates dependency parsing between words, pointwise mutual information (PMI) between regular expressions, cosine similarity between phrase embeddings, and KL divergence between topics in a heterogeneous graph for question and short document classification. The authors based the architecture on TensorGCN [10] and 4 heterogeneous node sets and adjacency matrices, while PQ-GCN is based on SHINE [11] and uses 5 homogeneous node sets and adjacency matrices which reduces sparsity in input data. The second work [2] uses word graph based on dependency parsing, and trains a graph attention network (GAT) [12]. This approach lacks consideration for other textual features, such as semantic and sequential relationships between words and phrases, that would enable better understanding of a question, given their short lengths.
II-B Text Classification Using Text Graphs
TextGCN [13] pioneered approaching text classification as a node classification problem. Yao et al. constructs a corpus-level graph with document nodes and word nodes—PMI as word-word edge weights and text frequency-inverse document frequency (TF-IDF) as word-document edge weights—and applies a 2-layer GCN to obtain the class output. Several works propose similar graph construction and node initialization based on TextGCN, with more efficient or better regularized graph propagation [14, 15].
Other works have also explored multi-graphs instead of a single corpus-level graph. TensorGCN [10] builds a tensor graph from three separate graphs: semantic graph via word embeddings, syntactic graph via grammar parsing, and sequential graph via PMI values. Each of these graphs has different word-word edge weights while sharing the same TF-IDF values for word-document edge weights, and they undergo intra- and inter-graph propagation. ME-GCN [16] constructs a graph that has multi-dimensional word-word, word-document, and document-document edges, using Word2Vec and Doc2Vec embeddings for node features, and normalized embedding distance and TF-IDF for edge features. SHINE [11] uses word graphs based on three word-level relationships—PMI, POS tag co-occurrence, and cosine similarities between embeddings of entity pairs appearing in the corpus—which are easy to obtain at a low cost using common natural language processing techniques.
While node classification is a transductive approach, where one needs to build a corpus-level graph consisting both train and test document nodes, InducT-GCN [17] extends the task into a inductive approach, where a training corpus graph is constructed first similarly to TextGCN, and any new test corpus is built as a virtual subgraph for inference.
Graph classification, in which you classify a document graph as opposed to a document node, has also been used as an inductive approach to text classification. Text-Level-GNN [18] uses a sliding window to build multiple graphs, each with a small number of nodes and edges, with trainable edge weights that are shared across the graphs for the same word pair. Huang et al. also uses a message passing mechanism instead of a GCN architecture, relying on aggregating neighborhood information via max-pooling and combining them via weighted sums, instead of convolution functions. Graph Fusion Network [19] constructs four individual global word co-occurrence graphs, and each document is represented as with a set of four subgraphs. GCN is applied to each subgraph, and the resulting output is concatenated, passed through an MLP layer, and average-pooled to obtain the final document embedding.
There are also combinatory methods based on GAT. HyperGAT (Hypergraph Attention Networks) [20] builds hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling. Attention mechanism is utilized at both node- and edge-level to update representations. Contrastive learning has also been utilized in conjunction with topic modeling and GAT to improve classification performance on short texts [21].
III Why Use Phrase Features?
As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions and yet be domain-agnostic to be generalizable across different classification frameworks. We identify a research gap in adapting GCNs for question classification in the lack of question classification-specific feature extraction and graph construction methods. While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations. The effectiveness of phrases in providing deeper semantic connections in the text has been discussed in topic modeling for questions [22, 1], as well as in graph-based document filtering tasks [23]. This section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
Figure 1: Corpus-level graph construction with extracted features
Capturing Meaningful Chunks of Information
Unlike individual words, phrases (e.g. noun phrases and verb phrases) often encapsulate more specific and contextually relevant information. Consider the question “How do enzymes function in the human body?” The phrase “enzymes function” and “human body” are pivotal in identifying this as a biological process question, likely classified under “Biology” or “Physiology.” The phrases indicates the action or process in question, and also specifies the domain of interest. In the question “Describe the steps to solve a quadratic equation,” the phrases “steps to solve” and “quadratic equation” are linked by their relevance to problem-solving in mathematics. This also applies to question analysis with part-of-speech (POS) tags. POS tags help in identifying the grammatical structure, but they do not capture the full semantic context when observed individually.
Intra-Question Relationships
By extracting phrases and analyzing their semantic embeddings, we can understand the relationship between different components of the question. For instance, in “What are the primary causes of the French Revolution?”, the proximity and relationship between “primary causes” and “French Revolution” are key to categorizing this question under “History” or “Political Science.”
Disambiguation and Contextualization
Phrases provide additional context that can help disambiguate words that might have multiple meanings. For instance, in “What is the role of the cell in an electric circuit?”, the phrase “role of the cell” and its proximity to “electric circuit” helps to clarify that “cell” refers to a component of an electric circuit rather than a biological cell.
While word-level features such as embeddings and POS tags are foundational for understanding the basic structure and meaning of a question, phrase-level features offer a more comprehensive and contextually rich understanding. By extracting noun phrases and verb phrases and analyzing their semantic embeddings and relationships, one can capture the nuanced meaning and intent of questions, leading to more accurate and effective classification models.
IV Methodology
IV-A Feature Extraction and Graph Construction
Given a corpus of question-label pair, we perform basic text cleaning which includes removal of contractions and punctuation.
To construct meaningful graph representations from a corpus of question text, we select several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data. A summary of the extracted features are provided in Table I.
TABLE I: Summary of Extracted Features
Node Type Node Feature Edge Weight
Word PPMI Word Word2Vec PMI value
Word POS POS tag - PMI value
Phrase PPMI Phrase PhraseBERT PMI value
Phrase POS Phrase POS tag - PMI value
Named Entities Named entity TransE Cosine similarity
Figure 2: Overview of model architecture
Words and PMI
We first tokenize the given question corpus by whitespace to obtain a unique set of words. A word-level graph is constructed, with word nodes connected by edges weighted with PMI values, representing local co-occurrence relationships. Specifically, we utilize positive PMI (PPMI) with a window size of 5, given by max(PMI($word_{1},word_{2}$), 0). PMI is given by:
$\text{PMI}(w_{1},w_{2})=\log\frac{P(w_{1},w_{2})}{P(w_{1})P(w_{2})}$ (1)
where $P(w_{1})=\frac{\#W(w_{1})}{\#W}$, $P(w_{2})=\frac{\#W(w_{2})}{\#W}$, $P(w_{1},w_{2})=\frac{\#W(w_{1},w_{2})}{\#W}$. #$\#W(w)$ is the number of sliding windows that a word $w$ in the corpus, and #$\#W(w_{1},w_{2})$ is the number of sliding windows that words $w_{1}$ and $w_{2}$ appear together in the corpus. #$\#W$ is the total number of sliding windows in the corpus.
Pre-trained word embeddings from word2vec [24] are also obtained for each word in the corpus, to be used as additional semantic information by concatenating with node embeddings.
TABLE II: PQ-GCN Model Parameters
Word Embedding Size 300
Phrase Embedding Size 768
Entity Embedding Size 100
Sliding Window Size for PMI 5
Hidden Dimension Size 200
No. of Layers in GCNs 2
Dropout 0.7
Threshold for Question Cosine Similarity 2.7
Optimizer Adam
Learning Rate 1e-3
Weight Decay 1e-4
Word-Level POS Tags
POS tags for each word is obtained via the default set from NLTK  [^1] https://nltk.org and PPMI values between each pair of POS tags are calculated as edge weights for a POS tag graph.
Phrases and PMI
Phrases are extracted from each question text via POS tag regex matching for noun and verb phrases with spaCy  [^2] https://spacy.io in Python, and a set of unique phrases in the corpus is obtained. The regex patterns for noun and verb phrases are given by (2) and (3) respectively:
[…]
$(\langle\text{NOUN}\rangle|\langle\text{PROPN}\rangle\langle\text{PART}\rangle?)+$ (2)
$\mu_{V}=$ $\langle\text{AUX}\rangle*\langle\text{ADV}\rangle*\langle\text{VERB}\rangle$ (3)
where ????, *, and $+$ represent zero or one, zero or more, and one or more occurrences of the preceding POS tag respectively.
Each question is represented as a sequence of verb phrases and noun phrases present in the text. Then a phrase-level graph (similar to the word-level graph) is constructed with PMI-weighted edges. Phrase embeddings are obtained from PhraseBERT [25] to provide additional semantic information.
TABLE III: Classification Label Distributions
Dataset NU ARC LREC Bloom TREC
Dataset Size 596 279 344 2522 5952
Classification Cognitive complexities Reasoning capabilities Expected answer types Educational objectives Question topics
Avg. # of Words 10.8 24.8 16.1 14.8 10.0
Class Labels Transfer (49.7%) Linguistic matching Very short answer (41.6%) Comprehension (38.1%) Entity (22.6%)
Knowledge (33.5%) (47.0%) Context sensitive (36.0%) Knowledge (13.6%) Description and abstract
Application (16.8%) Basic facts (28.0%) Answers will vary (22.4%) Evaluation (12.6%) concept (21.8%)
Hypothetical (25.0%) Application (12.5%) Human being (21.6%)
Analysis (12.1%) Numeric value (17.0%)
Synthesis (11.1%) Location (15.4%)
Abbreviation (1.6%)
TABLE IV: CNN Model Parameters
Parameter Layer 1 Layer 2 Layer 3 Layer 4
Type Conv2d Conv2d Conv2d Linear
Embedding Size 300
Sequence Length 15
Number of Filters 100 100 100
Kernel Size 3x3 4x4 5x5
Stride 1x1 1x1 1x1
Activation Function Softmax
Dropout Rate 0.5
TABLE V: Bi-LSTM Model Parameters
Parameter Layer 1 Layer 2 Layer 3
Type Bi-LSTM Bi-LSTM Linear
Embedding Size 300
Hidden Dimension Size 128 128
Activation Function Softmax
Phrase-Level POS Tags
[…]
Named Entity Recognition (NER)
Named entities (e.g., people, organizations, locations) carry significant meaning and often denote key elements within a question. A list of named entities is obtained from the NELL knowledge base [26], and is used to extract named entities via string matching. With the short text nature of questions, it is difficult to obtain multiple entities from a single question to calculate co-occurrence statistics. Hence, we utilize TransE embeddings [27] for each named entity and calculate the cosine similarity between embeddings of each pair of named entities, to be used as edge weights in constructing a named entity graph.
From the corpus, the above features are extracted and the respective relationships are computed to obtain edge weights, which are then used to construct the graph for each feature. The examples of graphs constructed are shown in Figure 1. For the corpus-level graphs constructed above, the node features are initialized as one-hot vectors except for entity graph, which uses the TransE embeddings directly as node features. The mapping of which text features are present in which question is used to generate question-level embeddings during propagation. By combining these diverse feature extraction methods, we construct a comprehensive and multi-faceted graph representation of the question text, capturing various semantic, syntactic, and named entity information crucial for question classification.
TABLE VI: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
Best F1, precision, and recall scores are in bold and the second best scores are underlined for each dataset.
Model Metric NU ARC LREC Bloom TREC
CNN F1 0.085 0.226 0.198 0.752 0.782
Precision 0.048 0.171 0.141 0.763 0.783
Recall 0.333 0.333 0.333 0.750 0.790
Bi-LSTM F1 0.607 0.564 0.480 0.425 0.653
Precision 0.606 0.554 0.511 0.465 0.692
Recall 0.620 0.598 0.495 0.422 0.642
TextGCN [13] F1 0.722 0.694 0.671 0.663 0.730
Precision 0.715 0.687 0.677 0.652 0.675
Recall 0.735 0.735 0.679 0.680 0.780
Text-Level-GNN [18] F1 0.185 0.200 0.404 0.287 0.623
Precision 0.694 0.362 0.424 0.250 0.651
Recall 0.183 0.275 0.424 0.344 0.606
HyperGAT [20] F1 0.715 0.372 0.745 0.173 0.678
Precision 0.733 0.386 0.766 0.176 0.673
Recall 0.703 0.374 0.737 0.178 0.691
TensorGCN [10] F1 0.412 0.499 0.566 0.107 0.805
Precision 0.421 0.557 0.692 0.147 0.851
Recall 0.450 0.604 0.575 0.157 0.775
SHINE [11] F1 0.560 0.610 0.620 0.459 0.560
Precision 0.553 0.600 0.620 0.461 0.568
Recall 0.583 0.623 0.627 0.473 0.613
ME-GCN [16] F1 0.632 0.603 0.601 0.607 0.659
Precision 0.738 0.601 0.607 0.629 0.718
Recall 0.619 0.620 0.608 0.592 0.634
InducT-GCN [17] F1 0.661 0.667 0.649 0.533 0.688
Precision 0.738 0.649 0.648 0.754 0.753
Recall 0.639 0.711 0.654 0.478 0.671
PQ-GCN F1 0.724 0.712 0.751 0.672 0.801
Precision 0.723 0.695 0.754 0.692 0.882
Recall 0.727 0.750 0.749 0.662 0.777
IV-B Model Architecture
As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as consolidating each graph of different node and edge types into a single question embedding. Initial approach was to concatenate the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function. Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
Following the initial exploration, we base our output layers largely on SHINE [11]. An overview of model architecture is provided in Figure 2. Each graph is passed through their own 2-layer GCN. Each GCN output is then concatenated to form question embeddings, and a question graph is created by using these question embeddings as node features and calculating the cosine similarity between the question embeddings for edge features. Hence, a dynamic graph of question nodes and cosine similarity edges is learned from the individual 2-layer GCN outputs, and this dynamic graph is propagated through a final 2-layer GCN and a linear layer for label classification. We provide the model and training parameters for the proposed PQ-GCN in Table II.
V Experiments
V-A Datasets
We first select a number of datasets used for question classification task to evaluate our proposed model on. The details are outlined in this subsection.
NU
This dataset was obtained from Najran University [28], containing 596 questions with 3 class labels of different cognitive complexities.
ARC
This dataset was released as part of the A12 Reasoning Challenge [29], consisting 279 questions labeled with 3 different reasoning capabilities.
LREC
This is a subset of 344 questions from the original dataset of science questions asked by teachers in real middle school classrooms, with a proposed taxonomy by expected answer types [7]. The subset was chosen by selecting 3 mutually exclusive class labels—’Very short answer’, ’Context sensitive’, and ’Answers will vary’.
Bloom
Bloom’s Taxonomy [30] is one of the most well-established question classification frameworks used by educators, based on different educational objectives. This dataset is a set of questions collected from different web sources and manually classified into the six cognitive levels of Bloom’s Taxonomy [^3] The dataset can be found at dx.doi.org/10.13140/RG.2.1.4932.3123.
TREC
The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC [31], about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set. These questions were manually labeled into 6 class labels.
Aside from TREC, for which the dataset split is obtained from HuggingFace  [^4] https://huggingface.co/datasets/CogComp/trec, the train/test splits are determined by a fixed seed and can be provided upon request. Relevant statistics for the datasets are found in Table III.
V-B Performance Evaluation
TABLE VII: Keyword Analysis for Five Datasets
Frequency is calculated by % of questions in the label set that contains the keyword.
Dataset Label Top Occurring Keywords (Frequency)
NU Transfer would (11.8%), compare (6.4%), story (6.4%)
Knowledge words (12.5%), describe (9.5%), state (7.0%)
Application construct (16%), use (14%), would (12.0%)
ARC Linguistic matching likely (22.1%), best (16.0%), water (10.7%)
Basic facts best(16.7%), chemical (12.8%), new (10.3%)
Hypothetical experiment (30.0%), water (27.1%), students (20.0%)
LREC Very short answer energy (13.3%), mass (11.9%), density (11.2%)
Context sensitive explain (18.6%), think (12.1%), lab (11.3%)
Answers will vary energy (28.6%), explain (20.8%), would (13.0%)
Bloom Comprehension explain (37.8%), discuss (16.6%), describe (14.3%)
Knowledge list (22.4%), define (16.0%), five (12.8%)
Evaluation justify (23.6%), evaluate (19.2%), answer (10.4%)
Application calculate (14.6%), explain (13.6%), determine (11.1%)
Analysis differentiate (17.8%), compare (15.5%), analyze (12.5%)
Synthesis suggest (23.7%), propose (11.83%), design (10.8%)
TREC Entity name (9.8%), fear (4.9%), first (3.9%)
Description and abstract concept mean (5.23%), origin (4.2%), get (2.8%)
Human being name (11.6%), first (7.38%), president (5.8%)
Numeric value many (32.7%), year (8.2%), much (6.1%)
Location country (13.7%), city (12.0%), state (7.2%)
Abbreviation stand (48.4%), abbreviation (19.0%), mean (11.6%)
We compare the performance of our proposed model against selected baseline models in Table VI, including vanilla CNN and Bi-LSTM for non-graph classification models.
The baseline models were selected by reviewing their contributions in graph neural networks for text classification [32], each using distinct feature engineering and graph construction methods for text data, as in Section II, as compared to works which have proposed architectural improvements for computational efficiency. With the exception of CNN and Bi-LSTM, for which parameters were set as outlined in Table IV and Table V, the rest of the baseline models have their model and training parameters set as provided by authors in the project repositories.
PQ-GCN achieves the best macro-averaged F1 scores across most datasets, including NU (0.724), LREC (0.751), and TREC (0.801), while consistently delivering high precision and recall. The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed in [11], showing the effectiveness of phrase-based features in improving question classification capabilities.
CNN model trained in this experiment is a relatively simple model with no specialized components, yet performs well on Bloom and TREC. This can be attributed to the characteristics of the two corpus, where presence of certain words indicate a strong association to a category. An analysis of most frequently occurring words have been included in Table VII, showing that the Bloom and TREC datasets contains distinct and mutually exclusive class-specific keywords. For example, Bloom’s Synthesis class contains high percentages of command words such as “suggest”, “propose”, and “design” that does not appear in questions of other classes. While the convolution kernels of CNN are effective in capturing these keywords as local features, it performs poorly on NU, ARC, and LREC where there are several overlapping non-class-specific keywords such as “would”, “water”, and “energy”.
TensorGCN performs with the highest F1 score on TREC. TensorGCN obtains its semantic embeddings by training an LSTM model on the fly during graph construction, and this method might be effective in generating more context-accurate embeddings with larger datasets. While this brings about a 0.004 point gain in F1 score over PQ-GCN, it also results in a longer graph construction time with increasing dataset size due to increased training time of the LSTM model, as compared to the proposed methodology of using pre-trained embedding models.
Bi-LSTM has a consistently average performance across the datasets, which is expected of a simple sequence model. TextGCN shows apt results for NU and ARC datasets where its F1 scores are second only to PQ-GCN, but there is room for improvement in leveraging text features other than PMI and TF-IDF when constructing the graph, especially when it comes to text with niche characteristics such as questions. Text-Level GNN shows inconsistent poor performance across the datasets, as its learnable word-word edge weights are highly dependent on relative proximity of keywords and the ability of the message passing mechanism to disambiguate polysemous words based on word proximity. HyperGAT also shows inconsistent performance, possibly attributed to its word-word edge features, which are extracted by segmenting a text at sentence level and treating each sentence as a “window”, as opposed to a sliding window in PMI calculation. Questions in the datasets usually consist of one or two sentences, and lacks usual word-sentence-document hierarchy that HyperGAT leverages on. ME-GCN’s embedding distance-based edge features and InducT-GCN’s text construction that mimics that of TextGCN also do not consider question-specific features, and PQGCN outperforms both models in F1 scores across all datasets.
VI Conclusion
We develop and evaluate PQ-GCN, adapting GCNs to question classification by incorporating phrase features into the text graphs that can provide deeper semantic connections within questions. The proposed method shows promising classification performance compared to baseline models, capturing the intricate word and phrase relationships within question text. Question classification is an important task in building educational analytics systems, as we move towards an era of personalized learning, augmented with artificial intelligence. Finding suitable feature extraction methods is crucial in tackling this task, and we hope to provide sufficient basis for further research into innovative feature extraction methods for automatic text classification with graph neural networks.